# Predict. Optimize. Revise. On Forecast and Policy Stability in Energy Management Systems

### v2

@@ II-C4 Accuracy of Determinstic Forecasts
II-C5 Accuracy of Probabilistic Forecasts
@@ III Experimentation and Results
According to the decision matrix in Figure I, at every time step the forecast and optimization plans can be updated or retained. Therefore, the forecast and optimization run with commitment periods $v_{F}$ and $v_{O}$. respectively. We evaluate the performance of the FHC algorithm with different combinations of forecast and optimization commitment. The result is shown in Tables II(a) and II(b). The scores in the tables indicate that updating the optimization plan more frequently than the forecast offers no added benefit. The optimal performance is achieved when the forecast and optimization are updated at the same frequency. Therefore, in the following analysis, we set $v=v_{F}=v_{O}$.
(a) FHC algorithm with deterministic forecast
@@
(b) FHC algorithm with stochastic forecast
@@
TABLE II: Optimization KPIs score for combinations of forecast and optimization commitment between 1 and 12
Further, we investigate the relationship between the performance of the FHC algorithm and properties of the forecast, namely accuracy and stability. In Figure 4, the accuracy and stability of the deterministic and stochastic forecasts are evaluated for different commitment periods. The comparison is made between the optimization with and without switching costs for the forecasts generated for out-of-sample buildings. Similar evaluation is done for the in-sample buildings and shown in Online Appendix D  [^4] Appendix D. [Online] Available: https://arxiv.org/abs/2407.03368. It is observed from the plots that the forecast error generally decreases with shorter commitment periods, while the vertical stability is better with longer commitment periods. The horizontal stability demonstrates a trade-off where the optimal level is reached at a certain commitment period between 1 and 12 hours.
For stochastic forecast error, measured with the Energy Score, we observe an anomaly when the lowest error is observed for scenarios updated every 8 hours. The hypothesis is that this is due to the nature of the implemented method of noise simulation. The magnitude of added gaussian noise is proportional to the value of the respective point prediction, which produces low variation in the regions around zero, where this specific configuration captures most accurately, therefore the resulting scenario sets capture the real distribution with a low variation.
The plots of the optimization scores show different behavior depending on the presence of the switching costs. The optimization score without switching costs is optimal with shorter commitment periods, while the addition of switching costs adds more fluctuation to this relationship. The optimal commitment period remains at 1 for the deterministic forecast, while the stochastic forecast enables an optimal performance at a longer commitment period.
These observations are supported by the correlation coefficients in the tables III(a) and III(b) with the correlation coefficients between the forecast properties and the optimization scores. The accuracy metrics are in positive correlation with the optimization score. When switching costs are factored in, the inverse correlation between the vertical stability and the cost becomes less pronounced. Yet, the correlation between the horizontal stability and the cost changes from negative to positive correlation when switching costs are included.
@@
Point Forecast + Deterministic MPC
Probabilistic forecast + Stochastic MPC
Figure 4: Accuracy (Top), Horizontal (Second) and Vertical Stability (Third row) of stochastic forecast over different revision periods with data from out-of-sample buildings. The optimization score is shown with and without switching costs. The forecast error is measured with MAE and Energy Score, while the stability is measured with MAC and EMD. The optimization score is measured with the cost function of the FHC algorithm.
(a) FHC algorithm with deterministic forecast
@@
(b) FHC algorithm with stochastic forecast
Correlation Coef. $ES$ $EMD_{V}$ $EMD_{H}$
$cost(FHC_{stochastic}(v))$ 0.63 -0.86 -0.29
$cost(FHC_{stochastic}(v))$ w/ switching costs 0.71 -0.06 0.47
TABLE III: Correlation Coefficients for Forecast Metrics and Optimization Score
IV Discussion
The observations from the results are consistent with the theoretical analysis of the upper bound on the competitive difference of the FHC algorithm. We observe that the optimal commitment period is dependent on the forecast properties. The inclusion of switching costs in the optimization problem changes the relationship between desired forecast properties and the optimization score. The forecast error is broadly lower with shorter commitment periods, however, the horizontal and vertical stability become also relevant. In the theoretical analysis in Online Appendix C, we see that for a more realistic model of the prediction noise with exponentially decaying correlation, a trade-off emerges between the forecast error and the switching costs. Through testing with real data, we observe that switching costs can be partially mitigated by committing to more stable forecasts and optimization plans.
@@ IV Discussion
Figure 5: The difference in net load between MPC with and without switching costs. Actual net load is shown in yellow. The base load is shown in green. The net load following the perfect forecast is shown in red. The electricity price is shown in blue.
@@
When prediction is inaccurate in the look-ahead window, the RHC is more effective than FHC with a commitment period. If we make predictions accurate for a small window $\gamma$ and then inaccurate for the remaining $(\omega-\gamma)$ steps of the lookahead window, the fixed horizon policy is affected by the inaccurate predictions whereas RHC only acts on the correct ones. This is typically the case for a multi-step horizon forecast when the nearest steps are predicted at a higher accuracy. The situation changes for a stochastic MPC, generally hedged better against prediction errors. We can see from the results that FHC with longer commitment period becomes viable when the optimization is computed for a set of scenarios, as opposed to a deterministic point forecast.
V Conclusion
In conclusion, this study highlights the critical role that switching costs, forecast accuracy, and stability play in the design and operation of energy management systems. In summary, our analysis shows that in systems where forecasts guide decision making, stability of predictions can reduce the frequency and severity of policy revisions, thereby mitigating the financial and operational impacts of switching costs. Our findings have significant implications for the design of energy management systems. They suggest that enhancing the stability of forecasts can improve system performance by reducing the need for frequent adjustments and managing the trade-offs between immediate and future resource allocation more effectively. Furthermore, we propose a novel approach to evaluating the stability of probabilistic forecasts using the Earth Mover’s Distance, which provides a robust measure of the dissimilarity between scenario sets.
@@ A-B MPC Formulation
  • $X_{t}^{MPC}(S_{t}|\theta^{LA})$ Represents the optimal set of decisions (battery charging $x_{t}^{pos}$ and discharing $x_{t}^{neg}$ actions) made by the MPC at time $t$ based on the current state $S_{t}$ and the lookahead approximation model $\theta^{LA}$ This model enables predicting future costs within a horizon $H$, taking into account the expected electric load and PV generation.
@@
A-C Theoretical Analysis of Performance Bounds
@@
A-D In-sample Performance
(a) Performance of the FHC algorithm with deterministic forecast. Study over combinations of forecast and optimization commitment between 1 and 12
@@ A-D In-sample Performance
(b) Performance of the FHC algorithm with stochastic forecast. Study over combinations of forecast and optimization commitment between 1 and 12
@@
TABLE IV: Optimization KPIs score for combinations of forecast and optimization commitment between 1 and 12 (with data from in-sample buildings)
@@
Point Forecast + Deterministic MPC
Probabilistic forecast + Stochastic MPC

### v3

@@ 2.3.4 Accuracy of Determinstic Forecasts
2.3.5 Accuracy of Probabilistic Forecasts
@@ 3 Experimentation and Results
According to the decision matrix in Figure 1, at every time step the forecast and optimization plans can be updated or retained. Therefore, the forecast and optimization run with commitment periods $v_{F}$ and $v_{O}$, respectively. We evaluate the performance of the FHC algorithm with different combinations of forecast and optimization commitment. The result is shown in Tables 2(a) and 2(b). The scores in the tables indicate that updating the optimization plan more frequently than the forecast offers no added benefit. The optimal performance is achieved when the forecast and optimization are updated at the same frequency. Therefore, in the following analysis, we set $v=v_{F}=v_{O}$.
(a) deterministic FHC algorithm with point forecast
@@
(b) stochastic FHC algorithm with probabilistic forecast
@@
Table 2: Optimization KPIs score for combinations of forecast and optimization commitment between 1 and 12
Further, we investigate the relationship between the performance of the FHC algorithm and properties of the forecast, namely accuracy and stability. In Figure 5, the accuracy and stability of the deterministic and stochastic forecasts are evaluated for different commitment periods. The comparison is made between the optimization with and without switching costs for the forecasts generated for out-of-sample buildings. Similar evaluation is done for the in-sample buildings and shown in A.4 It is observed from the plots that the forecast error generally decreases with shorter commitment periods, while the vertical stability is better with longer commitment periods. The horizontal stability demonstrates a trade-off where the optimal level is reached at a certain commitment period between 1 and 12 hours.
For stochastic forecast error, measured with the Energy Score, we observe an anomaly when the lowest error is observed for scenarios updated every 8 hours. The hypothesis is that this is due to the nature of the method of noise simulation, as outlined in [25]. The magnitude of added gaussian noise is proportional to the value of the respective point prediction, which produces low variation in the regions around zero, where this specific configuration captures most accurately, therefore the resulting scenario sets capture the real distribution with a low variation.
The plots of the optimization scores show different behavior depending on the presence of the switching costs. The optimization score without switching costs is optimal with shorter commitment periods, while the addition of switching costs adds more fluctuation to this relationship. The optimal commitment period remains at 1 for the point forecast, while the stochastic forecast enables an optimal performance at a longer commitment period.
These observations are supported by the correlation coefficients in the tables 3(a) and 3(b) with the correlation coefficients between the forecast properties and the optimization scores. The accuracy metrics are in positive correlation with the optimization score. When switching costs are factored in, the inverse correlation between the vertical stability and the cost becomes less pronounced. Yet, the correlation between the horizontal stability and the cost changes from negative to positive correlation when switching costs are included.
@@
Point Forecast
+ Deterministic FHC
Probabilistic forecast
+ Stochastic FHC
Figure 5: Accuracy (Top), Horizontal (Second) and Vertical Stability (Third row) of stochastic forecast over different revision periods with data from out-of-sample buildings.
(a) deterministic FHC algorithm
with point forecast
@@
(b) stochastic FHC algorithm
with probabilistic forecast
Correlation Coef. $ES$ $MASC_{V}$ $MASC_{H}$
$cost(FHC_{stoch.}(v))$ 0.63 -0.86 -0.29
$cost(FHC_{stoch.}(v))$ w/ switching costs 0.71 -0.06 0.47
Table 3: Correlation Coefficients for Forecast Metrics and Optimization Score
4 Discussion
The observations from the results are consistent with the theoretical analysis of the upper bound on the competitive difference of the FHC algorithm. We observe that the optimal commitment period is dependent on the forecast properties. The inclusion of switching costs in the optimization problem changes the relationship between desired forecast properties and the optimization score. The forecast error is broadly lower with shorter commitment periods, however, the horizontal and vertical stability become also relevant. In the theoretical analysis in A.3, we see that for a more realistic model of the prediction noise with exponentially decaying correlation, a trade-off emerges between the forecast error and the switching costs. Through testing with real data, we observe that switching costs can be partially mitigated by committing to more stable forecasts and optimization plans.
@@ 4 Discussion
Figure 6: The difference in net load between FHC with and without switching costs. Actual net load is shown in yellow. The base load is shown in green. The net load following the perfect forecast is shown in red. The electricity price is shown in blue.
@@
When prediction is inaccurate in the look-ahead window, the RHC is more effective than FHC with a commitment period. If we make predictions accurate for a small window $\gamma$ and then inaccurate for the remaining $(\omega-\gamma)$ steps of the lookahead window, the fixed horizon policy is affected by the inaccurate predictions whereas RHC only acts on the correct ones. This is typically the case for a multi-step horizon forecast when the nearest steps are predicted at a higher accuracy. The situation changes for a stochastic FHC, generally hedged better against prediction errors. We can see from the results that FHC with longer commitment period becomes viable when the optimization is computed for a set of scenarios, as opposed to a deterministic point forecast.
5 Conclusion
In conclusion, this study highlights the critical role that switching costs, forecast accuracy, and stability play in the design and operation of energy management systems. In summary, our analysis shows that in systems where forecasts guide decision making, stability of predictions can reduce the frequency and severity of policy revisions, thereby mitigating the financial and operational impacts of switching costs. Our findings have significant implications for the design of energy management systems. In the power system context, the switching costs are driven by imbalances, frequency deviations and additional stress on the power electronics. The current work suggests that enhancing the stability of forecasts can improve system performance by reducing the need for frequent adjustments and managing the trade-offs between immediate and future resource allocation more effectively. Furthermore, we propose $SDC$, a novel metric to evaluating the stability of probabilistic forecasts, which provides a robust measure of the dissimilarity between scenario sets.
@@ 5 Conclusion
Acknowledgments
This project has received funding from the European Union’s research and innovation programme Horizon Europe under the grant agreement No.101136211
@@ A.2 Optimization Formulation
  • 1.
    $X_{t}^{FHC}(S_{t}|\theta^{LA})$ Represents the optimal set of decisions (battery charging $x_{t}^{pos}$ and discharing $x_{t}^{neg}$ actions) made by the FHC at time $t$ based on the current state $S_{t}$ and the lookahead approximation model $\theta^{LA}$ This model enables predicting future costs within a horizon $H$, taking into account the expected electric load and PV generation.
  • 2.
@@
A.3 Theoretical Analysis of Performance Bounds
@@
A.4 In-sample Performance
(a) Performance of the FHC algorithm with point forecast.
Study over combinations of forecast and optimization commitment between 1 and 12
@@ A.4 In-sample Performance
(b) Performance of the FHC algorithm with stochastic forecast.
Study over combinations of forecast and optimization commitment between 1 and 12
@@
Table 4: Optimization KPIs score for combinations of forecast and optimization commitment between 1 and 12 (with data from in-sample buildings)
@@
Point Forecast + Deterministic FHC
Probabilistic forecast + Stochastic FHC